# Discrete Neural Flow Samplers with Locally Equivariant Transformer

### v1

@@ A.3 Connection to LEAPS
Equation 8 provides a natural foundation for learning the rate matrix $R_{t}^{\theta}$ using coordinate ascent. This involves alternating between two optimisation steps:
@@
A key insight is that even though Equation 8 formally holds when the expectation is taken under $p_{t}$, the training objective in Equation 18 remains valid for any reference distribution $q_{t}$ as long as it shares support with $p_{t}$. This is because, at optimality, the residual
@@
Integrating both sides with respect to $p_{t}$, and invoking the discrete Stein identity (Lemma 1), we find:
@@ Appendix B Derivation of Locally Equivariant Transformer
B.1 Proof of Proposition 1
It is worth noting that Proposition 1 was first introduced in (Zhang et al., 2023b, Proposition 5), and subsequently utilized by Campbell et al. (2024) to construct the conditional rate matrix, as well as by Holderrieth et al. (2025) in the development of the locally equivariant network. For completeness, we provide a detailed proof of Proposition 1 in this section.
@@ B.1 Proof of Proposition 1
We then restate Proposition 1 and provide a detailed proof as follows. See 1
@@ B.2 Locally Equivariant Networks
Based on Proposition 1, we can parametrise $R_{t}^{\theta}$ as a one-way rate matrix, which is theoretically capable of achieving the optimum that minimizes Equation 5. Although the one-way rate matrix is a restricted subset of general rate matrices and thus offers limited flexibility, it enables efficient computation of the objective in Equation 5. To see this, we first formally define the local equivariant network, originally proposed in Holderrieth et al. (2025).
@@
demonstrating the one-way property. With this parameterisation, the objective function in Equation 5 can be computed as
@@
where the second equality follows from the definition of the hollow network in Definition 3, which ensures that the $i$-th output is invariant to the changes in the $i$-th input. This confirms that $G$ satisfies the required local equivariance condition as in Definition 2
Based on Proposition 2, we present two locally equivariant architectures introduced in Holderrieth et al. (2025), followed by our proposed locally equivariant transformer.
@@ B.3 Locally Equivariant TransFormer (leTF)
In this section, we present the implementation details of the proposed Locally Equivariant Transformer (leTF), an expressive network architecture designed to preserve local equivariance. As introduced in Section 3.3, leTF is formulated as
@@
where $\omega$ denotes the learnable token embeddings produced by the projection layer (illustrated in Figure 2), and $H_{\mathrm{HTF}}$ represents the Hollow Transformer module. In the following, we focus on the implementation details of the hollow transformer $H_{\mathrm{HTF}}$.
Figure 8: Illustration of the hollow transformer.
As illustrated in Figure 8, the hollow transformer comprises $L$ bidirectional causal attention layers followed by a single attention readout layer. For clarity, we omit details of each causal attention layer, as they follow the standard Transformer architecture (Vaswani et al., 2017). We denote the outputs of the final left-to-right and right-to-left causal attention layers as
@@ C.1 Training and Sampling Algorithms
The training and sampling procedures are presented in Algorithms 1 and 2. For clarity, the rate matrix $R_{t}^{\theta}$ is parametrised using the proposed locally equivariant network, defined as $R_{t}^{\theta}(y,x)\triangleq[G_{t}^{\theta}(y_{i},i|x)]_{+}$, where $y$ and $x$ only differ at the $i$-th coordinate. To initiate training, we discretise the time interval $[0,1]$ into a set of evenly spaced time spans $\{t_{k}\}_{k=0}^{K}$, satisfying $10=t_{0}<\dots<t_{K}=$ 1 and $2t_{k}=t_{k+1}+t_{k-1}$ for all valid indices $k$.
In each outer loop of training, we generate trajectory samples by simulating the forward process under the current model parameters. Specifically, samples are drawn from the probability path $\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}^{\theta_{\mathrm{sg}}}}$, defined by the initial distribution $p_{0}$ and the current rate matrix $R_{t}^{\theta_{\mathrm{sg}}}$, where $\theta_{\mathrm{sg}}$ denotes $\mathrm{stop\_gradient}(\theta)$. This forward trajectory is simulated using the Euler–Maruyama method, as detailed in Algorithm 2, and stored in a replay buffer for reuse. During the inner training loop, we draw mini-batches uniformly from the buffer and compute the training loss based on Equation 10. The model parameters are then updated via gradient descent, as described in steps 6–11 of Algorithm 1.
@@
$\mathcal{B}\leftarrow\emptyset$triangleright Initialise replay buffer
@@
3:     $\{x_{{t}_{k}}^{(m)}\}_{m=1,k=0}^{M,K}\sim\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}^{\theta_{\mathrm{sg}}}}$triangleright Generate training samples
4:     $c_{t}\!\leftarrow\frac{1}{M}\sum_{m}\partial_{t}\log\tilde{p}_{t}(x_{t}^{(m)})-\sum_{y}R_{t}^{\theta_{\mathrm{sg}}}(x_{t}^{(m)},y)\frac{p_{t}(y)}{p_{t}(x_{t}^{(m)})},\forall t\in\{t_{k}\}_{k=0}^{K}$triangleright $\partial_{t}\log Z_{t}$
5:     $\mathcal{B}\leftarrow\mathcal{B}\cup\{(t_{k},x_{{t}_{k}}^{(m)})\}_{m=1,k=0}^{M,K}$triangleright update replay buffer
@@
7:         $\{(t,x_{t}^{(n)})\}_{n=1}^{N}\sim\mathcal{U}(\mathcal{B})$triangleright Uniformly sample from buffer
@@
9:         $\mathcal{L}(\theta)\leftarrow\frac{1}{N}\sum_{n}(\xi_{t}^{(n)}-c_{t})^{2}$triangleright Compute training loss
10:         $\theta\leftarrow\mathrm{optimizer\_step}(\theta,\nabla_{\theta}\mathcal{L}(\theta))$triangleright Perform gradient update
@@
1:$x_{0}\sim p_{0},\quad\Delta t\leftarrow\frac{1}{K},\quad t\leftarrow 0$triangleright Initialisation
@@
3:     $x_{t+\Delta t}\leftarrow\mathrm{Cat}(\mathbf{1}_{x_{t+\Delta t}=x}+R_{t}(x_{t+\Delta t},x)\Delta t)$triangleright Euler-Maruyama update
@@ C.2 Efficient Ratio Computation
where $\left[\log\frac{p_{t}(y)}{p_{t}(x)}\right]_{i,j}\triangleq\log\frac{p_{t}(x_{1},\dots,x_{i-1},j,x_{i+1},\dots,x_{d})}{p_{t}(x)}$ and we take the fact that $y$ and $x$ differ only in one position. This approximation requires computing $\nabla_{x}\log p_{t}(x)$ just once to estimate the ratio for the entire neighbourhood, thus improving computational efficiency. However, it introduces bias into the training objective in Equation 5, potentially leading to suboptimal solutions.
To illustrate this, we train DNFS to sample from pre-trained deep EBMs by minimising the objective in Equation 5 using two methods for computing the ratio $\frac{p_{t}(y)}{p_{t}(x)}$: i) exact computation in parallel; and ii) an approximation via the Taylor expansion in Equation 29. As shown Figure 9, using the approximate ratio leads to inaccurate samples that tend to lie in overly smoothed regions of the energy landscape. This degradation in sample quality is likely caused by the bias introduced by the approximation. Consequently, in deep EBM settings where the energy function is non-quadratic, we opt to compute the exact ratio in parallel, leaving the development of more efficient and unbiased approximations for future work.
Figure 9: Illustration of the effect of ratio computation in Equation 5. Top row: results using the Taylor expansion approximation; bottom row: results using exact computation. The approximated ratio yields less accurate samples, likely due to the bias introduced in the training objective.
@@ D.1 Importance Sampling
We begin by reviewing the fundamentals of importance sampling, followed by a proof based on the Radon-Nikodym derivative. This approach was also introduced in Lee et al. (2025); Holderrieth et al. (2025). For completeness, we include a brief recap to make the paper more self-contained.
@@
This estimator is consistent for any choice of $R_{t}$, and it becomes zero-variance when $R_{t}$ satisfies the Kolmogorov equation in Equation 4, a condition that can be approximately enforced by minimizing the loss in Equation 5. Before delving into the computation of the importance weights $\frac{\overleftarrow{\mathbb{Q}}^{\rho,R_{t}^{\prime}}}{\overrightarrow{\mathbb{Q}}^{\eta,R_{t}}}$, we introduce two key lemmas that underpin the derivation.
@@
where the first equation follows from Lemma 3, and the second from Lemma 4, leveraging the fact that $t\mapsto\log\tilde{p}_{t}$ is piecewise differentiable. The final equality holds by the detailed balance condition satisfied by the backward rate matrix, namely $R_{t}^{\prime}(y,x)p_{t}(x)=R_{t}(x,y)p_{t}(y)$. Importantly, although the partition function $Z_{t}$ is generally intractable, it cancels out in practice through the use of the self-normalised importance sampling estimator as in Equation 32.
Partition Function Estimation. To estimate the log-partition function $\log Z_{t}$, we consider the test function $\phi=\log\tilde{p}_{t}$, which yields the following lower bound:
$\log Z_{t}\geq{\mathbb{E}}_{p_{t}}[\log\tilde{p}_{t}]\!=\!{\mathbb{E}}_{x_{0\leq s\leq t}\sim\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}}}\!\left[\frac{\overleftarrow{\mathbb{Q}}^{p_{t},R_{t}^{\prime}}}{\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}}}\log\tilde{p}_{t}(x_{t})\right]\!\!\approx\!\sum_{k=1}^{K}\frac{\mathrm{exp}(w_{t}^{(k)})}{\sum_{j=1}^{K}\mathrm{exp}(w_{t}^{(j)})}\log\tilde{p}_{t}(x_{t}^{(k)}),$
@@
where the log-ratio $\log\frac{\overrightarrow{\mathbb{Q}}^{R_{t}^{\theta}|x_{0}}}{\overleftarrow{\mathbb{Q}}^{{R_{t}^{\theta}}^{\prime}|x_{1}}}$ can be evaluated using Lemma 3
@@ D.2 Training Discrete EBMs with Importance Sampling
where the second term can be easily approximated using the training data with Monte Carlo estimation. To estimate the intractable expectation over $p_{\phi}$, MCMC method is typically used. However, for computational efficiency, only a limited number of MCMC steps are performed, resulting in a biased maximum likelihood estimator and suboptimal energy function estimates (Nijkamp et al., 2020). To address this issue, we replace MCMC with the proposed discrete neural flow samplers. Specifically, we train a rate matrix $R_{t}^{\theta}$ to sample from the target EBM $p_{\phi}$. The expectation over $p_{\phi}$can then be estimated using CTMT-inspired importance sampling, as described in Section D.1:
@@
    Updating the rate matrix parameters $\theta$ using the training procedure described in Algorithm 1;
@@
    Updating EBM $p_{\phi}$ via contrastive divergence, as defined in Equation 44.

### v2

@@ A.3 Connection to LEAPS
Equation˜8 provides a natural foundation for learning the rate matrix $R_{t}^{\theta}$ using coordinate ascent. This involves alternating between two optimisation steps:
@@
A key insight is that even though Equation˜8 formally holds when the expectation is taken under $p_{t}$, the training objective in Equation˜18 remains valid for any reference distribution $q_{t}$ as long as it shares support with $p_{t}$. This is because, at optimality, the residual
@@
Integrating both sides with respect to $p_{t}$, and invoking the discrete Stein identity (Lemma˜1), we find:
@@
Since NDFS closely resembles LEAPS, we summarise the main distinctions below to highlight the unique contributions of our work:
  • While LEAPS and DNFS yield similar objective functions, they are derived from different perspectives. DNFS derives the objective by learning the rate matrix to satisfy the Kolmogrove equation, whereas LEAPS learns the rate matrix by minimising the importance weights. Perhaps surprisingly, these two perspectives lead to similar objectives. However, the new perspective from the Kolmogorov equation offers a new insight for future research: leveraging more accurate estimators of $\partial_{t}\log Z_{t}$ to further improve performance, which is not evident from the LEAPS framework.
  • The success of DNFS highly depends on the proposed Locally Equivariant Transformer (leTF). Compared to the locally equivariant networks in LEAPS, leTF offers greater model capacity and improved adaptability, making it more suitable for diverse modalities and complex input structures. We hope this architectural advancement will inspire future developments, which are essential for advancing both LEAPS and DNFS.
  • Unlike LEAPS, which is only evaluated on synthetic Ising and Potts models, DNFS is tested on broader applications, including sampling from Ising models, training EBMs, and solving combinatorial optimisation problems. We hope this wider empirical scope will inspire further research into additional applications of discrete neural samplers.
@@ Appendix B Derivation of Locally Equivariant Transformer
B.1 Proof of Proposition˜1
It is worth noting that Proposition˜1 was first introduced in (Zhang et al., 2023b, Proposition 5), and subsequently utilized by Campbell et al. (2024) to construct the conditional rate matrix, as well as by Holderrieth et al. (2025) in the development of the locally equivariant network. For completeness, we provide a detailed proof of Proposition˜1 in this section.
@@ B.1 Proof of Proposition˜1
We then restate Proposition˜1 and provide a detailed proof as follows. See 1
@@ B.2 Locally Equivariant Networks
Based on Proposition˜1, we can parametrise $R_{t}^{\theta}$ as a one-way rate matrix, which is theoretically capable of achieving the optimum that minimizes Equation˜5. Although the one-way rate matrix is a restricted subset of general rate matrices and thus offers limited flexibility, it enables efficient computation of the objective in Equation˜5. To see this, we first formally define the local equivariant network, originally proposed in Holderrieth et al. (2025).
@@
demonstrating the one-way property. With this parameterisation, the objective function in Equation˜5 can be computed as
@@
where the second equality follows from the definition of the hollow network in Definition˜3, which ensures that the $i$-th output is invariant to the changes in the $i$-th input. This confirms that $G$ satisfies the required local equivariance condition as in Definition˜2
Based on Proposition˜2, we present two locally equivariant architectures introduced in Holderrieth et al. (2025), followed by our proposed locally equivariant transformer.
@@ B.3 Locally Equivariant TransFormer (leTF)
Figure 8: Illustration of the hollow transformer.
In this section, we present the implementation details of the proposed Locally Equivariant Transformer (leTF), an expressive network architecture designed to preserve local equivariance. As introduced in Section˜3.3, leTF is formulated as
@@
where $\omega$ denotes the learnable token embeddings produced by the projection layer (illustrated in Figure˜2), and $H_{\mathrm{HTF}}$ represents the Hollow Transformer module. In the following, we focus on the implementation details of the hollow transformer $H_{\mathrm{HTF}}$.
As illustrated in Figure˜8, the hollow transformer comprises $L$ bidirectional causal attention layers followed by a single attention readout layer. For clarity, we omit details of each causal attention layer, as they follow the standard Transformer architecture (Vaswani et al., 2017). We denote the outputs of the final left-to-right and right-to-left causal attention layers as
@@ C.1 Training and Sampling Algorithms
The training and sampling procedures are presented in Algorithms˜1 and 2. For clarity, the rate matrix $R_{t}^{\theta}$ is parametrised using the proposed locally equivariant network, defined as $R_{t}^{\theta}(y,x)\triangleq[G_{t}^{\theta}(y_{i},i|x)]_{+}$, where $y$ and $x$ only differ at the $i$-th coordinate. To initiate training, we discretise the time interval $[0,1]$ into a set of evenly spaced time spans $\{t_{k}\}_{k=0}^{K}$, satisfying $0=t_{0}<\dots<t_{K}=1$ and $2t_{k}=t_{k+1}+t_{k-1}$ for all valid indices $k$.
In each outer loop of training, we generate trajectory samples by simulating the forward process under the current model parameters. Specifically, samples are drawn from the probability path $\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}^{\theta_{\mathrm{sg}}}}$, defined by the initial distribution $p_{0}$ and the current rate matrix $R_{t}^{\theta_{\mathrm{sg}}}$, where $\theta_{\mathrm{sg}}$ denotes $\mathrm{stop\_gradient}(\theta)$. This forward trajectory is simulated using the Euler–Maruyama method, as detailed in Algorithm˜2, and stored in a replay buffer for reuse. During the inner training loop, we draw mini-batches uniformly from the buffer and compute the training loss based on Equation˜10. The model parameters are then updated via gradient descent, as described in steps 6–11 of Algorithm˜1.
@@
[…]
@@
3:  $\{x_{{t}_{k}}^{(m)}\}_{m=1,k=0}^{M,K}\sim\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}^{\theta_{\mathrm{sg}}}}$ $\triangleright$ Generate training samples
4:  $c_{t}\!\leftarrow\frac{1}{M}\sum_{m}\partial_{t}\log\tilde{p}_{t}(x_{t}^{(m)})-\sum_{y}R_{t}^{\theta_{\mathrm{sg}}}(x_{t}^{(m)},y)\frac{p_{t}(y)}{p_{t}(x_{t}^{(m)})},\forall t\in\{t_{k}\}_{k=0}^{K}$ $\triangleright$ $\partial_{t}\log Z_{t}$
5:  $\mathcal{B}\leftarrow\mathcal{B}\cup\{(t_{k},x_{{t}_{k}}^{(m)})\}_{m=1,k=0}^{M,K}$ $\triangleright$ update replay buffer
@@
7:   $\{(t,x_{t}^{(n)})\}_{n=1}^{N}\sim\mathcal{U}(\mathcal{B})$ $\triangleright$ Uniformly sample from buffer
@@
9:   $\mathcal{L}(\theta)\leftarrow\frac{1}{N}\sum_{n}(\xi_{t}^{(n)}-c_{t})^{2}$ $\triangleright$ Compute training loss
10:   $\theta\leftarrow\mathrm{optimizer\_step}(\theta,\nabla_{\theta}\mathcal{L}(\theta))$ $\triangleright$ Perform gradient update
@@
[…]
@@
3:  $x_{t+\Delta t}\leftarrow\mathrm{Cat}(\mathbf{1}_{x_{t+\Delta t}=x}+R_{t}(x_{t+\Delta t},x)\Delta t)$ $\triangleright$ Euler-Maruyama update
@@ C.2 Efficient Ratio Computation
where $\left[\log\frac{p_{t}(y)}{p_{t}(x)}\right]_{i,j}\triangleq\log\frac{p_{t}(x_{1},\dots,x_{i-1},j,x_{i+1},\dots,x_{d})}{p_{t}(x)}$ and we take the fact that $y$ and $x$ differ only in one position. This approximation requires computing $\nabla_{x}\log p_{t}(x)$ just once to estimate the ratio for the entire neighbourhood, thus improving computational efficiency. However, it introduces bias into the training objective in Equation˜5, potentially leading to suboptimal solutions.
To illustrate this, we train DNFS to sample from pre-trained deep EBMs by minimising the objective in Equation˜5 using two methods for computing the ratio $\frac{p_{t}(y)}{p_{t}(x)}$: i) exact computation in parallel; and ii) an approximation via the Taylor expansion in Equation˜29. As shown Figure˜9, using the approximate ratio leads to inaccurate samples that tend to lie in overly smoothed regions of the energy landscape. This degradation in sample quality is likely caused by the bias introduced by the approximation. Consequently, in deep EBM settings where the energy function is non-quadratic, we opt to compute the exact ratio in parallel, leaving the development of more efficient and unbiased approximations for future work.
Figure 9: Illustration of the effect of ratio computation in Equation˜5. Top row: results using the Taylor expansion approximation; bottom row: results using exact computation. The approximated ratio yields less accurate samples, likely due to the bias introduced in the training objective.
@@ D.1 Importance Sampling
We begin by reviewing the fundamentals of importance sampling, followed by a proof based on the Radon-Nikodym derivative. This approach was also introduced in Lee et al. (2025); Holderrieth et al. (2025); Pani et al. (2025). For completeness, we include a brief recap to make the paper more self-contained.
@@
This estimator is consistent for any choice of $R_{t}$, and it becomes zero-variance when $R_{t}$ satisfies the Kolmogorov equation in Equation˜4, a condition that can be approximately enforced by minimizing the loss in Equation˜5. Before delving into the computation of the importance weights $\frac{\overleftarrow{\mathbb{Q}}^{\rho,R_{t}^{\prime}}}{\overrightarrow{\mathbb{Q}}^{\eta,R_{t}}}$, we introduce two key lemmas that underpin the derivation.
@@
where the first equation follows from Lemma˜3, and the second from Lemma˜4, leveraging the fact that $t\mapsto\log\tilde{p}_{t}$ is piecewise differentiable. The final equality holds by the detailed balance condition satisfied by the backward rate matrix, namely $R_{t}^{\prime}(y,x)p_{t}(x)=R_{t}(x,y)p_{t}(y)$. Importantly, although the partition function $Z_{t}$ is generally intractable, it cancels out in practice through the use of the self-normalised importance sampling estimator as in Equation˜32.
Free Energy and Internal Energy Estimation. To estimate the log-partition function $\log Z_{t}$, we consider the following lower bound
$\log Z_{t}$ $=\log{\mathbb{E}}_{p_{t}}\left[\frac{Z_{t}}{Z_{0}}\right]=\log{\mathbb{E}}_{x_{0\leq s\leq t}\sim\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}}}\left[\frac{\overleftarrow{\mathbb{Q}}^{p_{t},R_{t}^{\prime}}}{\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}}}\frac{Z_{t}}{Z_{0}}\right]$
$\geq{\mathbb{E}}_{x_{0\leq s\leq t}\sim\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}}}\left[\log\frac{\overleftarrow{\mathbb{Q}}^{p_{t},R_{t}^{\prime}}}{\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}}}+\log\frac{Z_{t}}{Z_{0}}\right]$
$={\mathbb{E}}_{x_{0\leq s\leq t}\sim\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}}}\left[\int_{0}^{t}\partial_{s}\log\tilde{p}_{s}(x_{s})-\sum_{y}R_{s}(x_{s},y)\frac{p_{s}(y)}{p_{s}(x_{s})}\mathop{}\!\mathrm{d}s\right],$ (37)
where we assume that $Z_{0}=1$. Considering the test function $\phi=\log\tilde{p}_{t}$, one can also estimate the negative entropy, which is related to the internal energy:
${\mathbb{E}}_{p_{t}}[\log\tilde{p}_{t}]\!=\!{\mathbb{E}}_{x_{0\leq s\leq t}\sim\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}}}\!\left[\frac{\overleftarrow{\mathbb{Q}}^{p_{t},R_{t}^{\prime}}}{\overrightarrow{\mathbb{Q}}^{p_{0},R_{t}}}\log\tilde{p}_{t}(x_{t})\right]\!\!\approx\!\sum_{k=1}^{K}\frac{\mathrm{exp}(w_{t}^{(k)})}{\sum_{j=1}^{K}\mathrm{exp}(w_{t}^{(j)})}\log\tilde{p}_{t}(x_{t}^{(k)}),$ (38)
@@
where the log-ratio $\log\frac{\overrightarrow{\mathbb{Q}}^{R_{t}^{\theta}|x_{0}}}{\overleftarrow{\mathbb{Q}}^{{R_{t}^{\theta}}^{\prime}|x_{1}}}$ can be evaluated using Lemma˜3
@@ D.2 Training Discrete EBMs with Importance Sampling
where the second term can be easily approximated using the training data with Monte Carlo estimation. To estimate the intractable expectation over $p_{\phi}$, MCMC method is typically used. However, for computational efficiency, only a limited number of MCMC steps are performed, resulting in a biased maximum likelihood estimator and suboptimal energy function estimates (Nijkamp et al., 2020). To address this issue, we replace MCMC with the proposed discrete neural flow samplers. Specifically, we train a rate matrix $R_{t}^{\theta}$ to sample from the target EBM $p_{\phi}$. The expectation over $p_{\phi}$can then be estimated using CTMT-inspired importance sampling, as described in Section˜D.1:
@@
    Updating the rate matrix parameters $\theta$ using the training procedure described in Algorithm˜1;
@@
    Updating EBM $p_{\phi}$ via contrastive divergence, as defined in Equation˜46.